Write an equation of the line containing the specified point and perpendicular to the indicated line .
(0,-7), 4x+3=y
Thank you
Jennifer

Answer by lwsshak3(11628)

Write an equation of the line containing the specified point and perpendicular to the indicated line .
(0,-7), 4x+3=y
***
equation of a line: y=mx+b, m=slope, b=y-intercept
given line:4x+3=y
y=4x+3
slope=4
..
slope of a line perpendicular to given line=-1/4 (negative reciprocal)
equation:
y=-x/4+b
solving for b using coordinates of given point (0,-7)
-7=-0/4+b
b=-7
equation of perpendicular line:
y=-x/4-7